### v1

@@ 2.1 Upper Bounds for Stochastic Processes
Once the expected value is understood, tail bounds can be established either by applying concentration of measure, or by reworking the chaining steps to control the tail probability instead of the expectation. For technical convenience, our discussion of tail bounds will be restricted to processes of the form $X_{t}=\left<Y,t\right$>, where $t\in T\subseteq\ell_{2}$. Following Vershynin, (2018), the sub-Gaussian norm of a random vector $Y\in\ell_{2}$ is defined by taking the supremum over one-dimensional projections:
$\|Y\|_{\Psi_{2}}:=\sup_{v\in\ell_{2}}\|\left<Y,v\right>\|_{\Psi_{2}}.$ (16)
A bound on $\|Y\|_{\Psi_{2}}$ can be guaranteed, for example, if $Y$ has independent coordinates with uniformly bounded sub-Gaussian norms (Vershynin,, 2018).
Concentration inequalities control the deviation probability from the mean. We say $P_{Y}$ satisfies the log-Sobolev inequality (LSI), if for any continuously differentiable function $f\colon\ell_{2}\to\mathbb{R}$,
${\rm Ent}(f^{2})\leq 2\mathbb{E}[\|\nabla f(Y)\|_{2}^{2}],$ (17)
where ${\rm Ent}(f):=\mathbb{E}[f(Y)\ln f(Y)]-\mathbb{E}[f(Y)]\ln\mathbb{E}[f(Y)]$. The log-Sobolev inequality is satisfied when the coordinates of $Y$ are independent and each satisfies the log-Sobolev inequality (tensorization property). The log-Sobolev inequality implies sub-Gaussian concentration. In fact, the following property is well-known:
Lemma 1.
Suppose that $X_{t}:=\left<t,Y\right$>, where $Y$ is a zero-mean random vector in $\ell_{2}$ satisfying the log-Sobolev inequality, and $T$ is totally bounded (i.e., covering numbers are all finite) in $\ell_{2}$. Let $\sigma^{2}:=\sup_{t\in T}\mathbb{E}[X_{t}^{2}]$, and let $\xi:=\sup_{t\in T}X_{t}$. Then for all $u>0$,
$\mathbb{P}[\xi-\mathbb{E}[\xi]>u]\leq e^{-\frac{u^{2}}{2\sigma^{2}}}.$ (18)
The proof follows by showing the Lipschitz condition $\|\nabla\max_{t\in T}\left<t,y\right>\|_{2}\leq\max_{t\in T}\|t\|_{2}$, and using the fact that the log-Sobolev inequality implies sub-Gaussian concentration; see (Boucheron et al.,, 2013, Theorem 5.8). In fact, from the proof we can see that the LSI assumption in Lemma 1 can be weakened to Talagrand’s optimal transport inequality, which also implies sub-Gaussian deviation of Lipschitz functions (Boucheron et al.,, 2013).
It is possible, however, to obtain tail bounds without assuming concentration:
Lemma 2.
Suppose that $Y$ is a zero-mean random vector in $\ell_{2}$ satisfying $\|Y\|_{\Psi_{2}}\leq 1$, and $\sup_{t\in T}\|t\|_{2}\leq\Delta$. Then there exists a universal constant $C>0$ such that for any $x\geq 0$,
$\mathbb{P}[\sup_{t\in T}\left<t,Y\right>\geq CJ(T)+x]\leq Ce^{-\frac{x^{2}}{C\Delta^{2}}},$ (19)
where $J(T):=\int_{0}^{\infty}\sqrt{\ln N(T,\epsilon)}\epsilon$.
A proof of Lemma 2 can be found in Van Handel, (2014) and the references therein. Statistical literature, such as Van de Geer, (1990), also used similar tail probability bounds for sub-Gaussian processes. As noted in Van Handel, (2014), although Lemma 2 requires weaker an assumption on the distribution than Lemma 1, it can give a weaker bound when $J(T)$ is larger than $\mathbb{E}[\xi]$.
@@ 2.2 Empirical Risk Minimization (ERM)
In a general mean estimation problem, $T$ is a closed subset of $\ell_{2}$, and $X_{1},\dots,X_{n}$ are i.i.d. random vectors in $\ell_{2}$ whose mean is an unknown $m\in T$. For any $t\in\ell_{2}$, define the empirical risk
@@
where $\chi_{z}:=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, and $(\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric (up to a universal constant factor). An example of such a mean estimation problem, which plays a fundamental role in nonparametric statistics, will be discussed in Section 3.4, where $T$ is a Sobolev ellipsoid.
In a general regression problem, $T$ is a closed subset in $\mathbb{R}^{n}$, and $Y=\sqrt{n}m+w$, where $m\in T$ is an unknown vector to estimate, and $w$ is a zero-mean noise vector satisfying $\|w\|_{\Psi_{2}}\leq 1$. An example of sparse regression with fixed designs will be discussed in Section 3.5. The case of random designs, which is traditionally tackled by local Rademacher complexity techniques (Lugosi and Wegkamp,, 2004; Bartlett et al.,, 2005), is not considered here, although it would be interesting to consider such extensions for future work. Define
$\hat{L}(t):=\frac{1}{n}\|\sqrt{n}t-Y\|_{2}^{2}$ (25)
and define $L(t)$ and $\hat{m}$ the same way as in the mean estimation problem. Then (24) still holds, if we now define $\chi_{z}:=-2\left<z,w\right$>, and $(\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric (up to a universal constant factor).
With a slight change of notation, it is possible to unify the formulations of the two problems: both are essentially about projecting a random vector onto a subset in $\ell_{2}$. We present them as estimation and regression problems separately simply to make the sample size $n$ emerge naturally and consistently with the nonparametric estimation and sparse regression literature. More generally, the framework also encompasses many other problems including functional estimation, matrix completion, shape-restricted regression, and isotonic regression; see Chatterjee, (2014); Chatterjee et al., (2015).
@@
In both problems, the inequality (24) plays a key role in upper bounding the mean square error $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]$; see e.g. (Van de Geer,, 1990, eq (2)). Since $\mathbb{E}[\hat{L}(m)-L(m)]=0$, the expected generalization error equals $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], which is also connected to (24). If we take the expectation of both sides of (24), and upper bound the right side by the Dudley integral using the covering number of $T$, we generally do not get sharp convergence rates of the mean square error. In fact, even for the simplest case of 1-dimensional parameter estimation, such a naive approach yields a $1/\sqrt{n}$ rate rather than the optimal $1/n$ rate.
@@
The traditional approach for improving the bound is a “slicing” argument that yields local complexity bounds; see for example Van de Geer, (1990); Wainwright, (2019); Raskutti et al., (2011) or (Van Handel,, 2014, Theorem 5.32). Roughly speaking, if $\|\hat{m}-m\|^{2}=r$, then $\chi_{m-\hat{m}}\leq\sup_{z\in B(r)\cap(T-T)}\chi_{z}$, which is bounded by some function $\omega(r)$ with high probability. The function $\omega(r)$ typically grows slower than $r^{2}$, so $r^{2}\lesssim\frac{1}{\sqrt{n}}\omega(r)$ yields an upper bound on $r$. More precisely, a union bound argument is applied to control the generalization error at different scales of $r$ with high probability. While it may seem that the uniform control by the union bound could lose logarithmic factors, this can actually be avoided by carefully scheduling of the error tolerance at each scale.
@@
In Section 8.1, we give a more detailed review of how the traditional slicing approach can yield sharp convergence for the one-dimensional toy problem. Local complexity bounds, which are based on Dudley’ integral for local sets, can still be loose for general $T$ where Dudley’ integral is loose. The majorizing measure theorem, though sharp, is hard to evaluate, and is rarely used in the analysis of ERM.
@@ 2.3 Information-Theoretic Tools
We recall standard information-theoretic measures in Thomas and Joy, (2006), using the natural base of the logarithm. The entropy of a discrete random variable $X\sim\mu$ is $H(X):=\sum_{x}\mu(x)\ln\frac{1}{\mu(x)}$. The relative entropy between $\mu$ and $\nu$ on the same measurable space is $D(\mu\|\nu):=\int\ln\frac{d\mu}{d\nu}d\mu$. Mutual information between $(X,Y)\sim P_{XY}$ is defined as $I(X;Y):=D(P_{XY}\|P_{X}\times P_{Y})$. Conditional versions of these quantities are defined by averaging over the random variable which is conditioned on. For example, $D(P_{\hat{Z}|Z}\|\mu|P_{Z}):=\int D(P_{\hat{Z}|Z=z}\|\mu)dP_{Z}(z)$. The lifting argument relies on the method of types, a standard method in information theory (Csiszár and Körner,, 2011) and large deviation analysis (Dembo,, 2009). The type of a sequence of length $N$ is simply its empirical distribution. Since not all distributions can be a type, we need the following:
@@
Suppose that $\mu$ is a distribution on a finite set $\mathcal{Z}$. We say $\mu$ is rational if there exists integer $N>0$ such that $N\mu$ is integer, i.e., $N\mu(z)$ is integer for any $z\in\mathcal{Z}$.
If $N\nu$ is integer, we denote by $C_{\nu}$ the $\nu$-type set, i.e., the set of sequences with type $\nu$. Then the following result is standard and can be found in (Csiszár and Körner,, 2011, Chapter 1):
@@
Let $S_{N}$ be the set of all possible types of $\mathcal{Z}^{N}$, where $\mathcal{Z}$ is a set satisfying the cardinality bound $|\mathcal{Z}|\leq n$. Then $|S_{N}|={n+N-1\choose n-1}\leq(N+1)^{n}$. Furthermore, if $\mu$ is a distribution on $\mathcal{Z}$ and $N\mu$ is integer, then
$(N+1)^{-n}\exp(-ND(\nu\|\mu))\leq\mu^{\otimes N}(C_{\nu})\leq\exp(-ND(\nu\|\mu)).$ (26)
@@ 3.1 A Rate-Distortion Integral
with ties broken arbitrarily but respecting measurability. Assuming measurability of $Z(\omega)$, we have measurability of $d(Z(\omega),t_{j})$ for each $j$, so that $Z_{k}$, and hence $X_{Z_{k}}$, is measurable. Density of $T_{0}$ ensures $\lim_{k\to\infty}Z_{k}=Z$ almost surely, and continuity implies that $X_{Z}=\lim_{k\to\infty}X_{Z_{k}}$ is measurable, being a sequential limit of measurable functions.
We recall the rate-distortion function in information theory (Thomas and Joy,, 2006), which will be a counterpart of the covering number in Dudley’s integral:
@@
Given $\mu$ on a metric space $(T,d)$, the rate-distortion function $R_{\mu}(\sigma^{2}):=\inf_{P_{UZ}}I(U;Z)$, where the infimum is over $P_{UZ}$ satisfying $P_{Z}=\mu$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$.
Given a probability measure $\mu$ on $T$ and $\sigma>0$, define
Then we have
$R_{\mu}(\sigma^{2})\leq\inf_{P_{Z\hat{Z}}\in\Pi_{\sigma}(P_{Z})}I(Z;\hat{Z})\leq R_{\mu}(\sigma^{2}/4).$ (30)
[…]
For $Z\sim\mu$, define
$\sigma_{\rm m}(\mu):=\inf_{z\in T}\sqrt{\mathbb{E}[d^{2}(Z,t)]},$ (31)
Definition 3.
Given a stochastic process $(X_{t}(\omega))_{t\in T}$, define the complexity
@@
If $(X_{t})_{t\in T}$ is a centered sub-Gaussian process on a countable set $T$, and $\mu$ is a distribution on $T$ for which $\sigma_{\rm m}<\infty$, then
$\mathsf{w}(\mu)\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma.$ (33)
The proof is given in Section 4. It will be seen in Lemma 6 that if the right side of (33) is finite, then $\mathbb{E}[|X_{Z}|]<\infty$, so no need to worry about absolute integrability on the left side of (33). The constant 48 in the theorem is from the commonly referenced estimate $K<24$ in the Dudley integral (7) for Gaussian processes (see comments after (7) for the extension to the sub-Gaussian case). Actually, according to Dudley, (2016), the best published estimate appears to be $K\leq 4\sqrt{2}$ given in Lifshits, (2012), which would imply that 48 in Theorem 1 can be improved to $8\sqrt{2}$.
@@
Suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, and let $Y:=(X_{t})\in\mathbb{R}^{|T|}$. Let $Z$ be a random variable on $T$ arbitrarily correlated with $Y$, and set $\mu:=P_{Z}$. Then
$\mathbb{E}[X_{Z}]\leq 48\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})\wedge I(Y;Z)}d\sigma.$ (34)
The proof of Theorem 2 is given in Section 4.1. Since $\sigma_{\rm m}\lesssim\sup_{t\in T}\|X_{t}\|_{\Psi_{2}}$, it is easy to see that Theorem 2 implies (2). Since the rest of the paper does not concern Theorem 2, we will not discuss its extension beyond finite $T$. While Theorem 1 does not cover the case of uncountable $T$, it is sufficient for our statistical applications due to the following observation:
@@
If $(X_{t})$ is a separable, continuous process, then for $Z\sim\mu$,
$\mathbb{E}[X_{Z}]\leq 48\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(\mu,\mu^{\prime})\leq\epsilon}\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (35)
The proof of Proposition 4.2 is in Section 4.2. Naturally, we would like to show that the right side of (35) reduces to the right side of (33) by dominated convergence, but this requires showing that $\sqrt{R_{\mu^{\prime}}(\sigma^{2})}$ can be dominated by an absolutely integrable function, which is challenging in the regime of vanishing $\sigma$ where the rate-distortion function may grow fast. While it is possible to impose on certain growth condition for dominated convergence to apply, we will not dive into those complications. Instead, in many statistical applications, we will see that $R_{\mu^{\prime}}(\sigma^{2})$ can controlled using the second moments of $\mu$, so the right side of (35) can be directly evaluated using the continuity of the bound in the second moment, without the need to establish the equivalence of the right sides of (35) and (33).

### v3

@@ 2.2 Empirical risk minimization (ERM)
Consider the mean estimation problem, where $T$ is a closed subset of $\ell_{2}$, and $X_{1},\dots,X_{n}$ are i.i.d. random vectors in $\ell_{2}$ whose mean is an unknown $m\in T$. We assume that $\|X_{1}-m\|_{G}\leq 1$, where following [43], the sub-Gaussian norm of a random vector $Y\in\ell_{2}$ is defined by taking the supremum over one-dimensional projections:
$\|Y\|_{G}:=\sup_{v\colon\|v\|_{2}=1}\|\left<Y,v\right>\|_{G}.$ (13)
A bound on $\|Y\|_{G}$ can be guaranteed, for example, if $Y$ has independent coordinates with uniformly bounded sub-Gaussian norms [43]. For any $t\in\ell_{2}$, define the empirical risk
@@
where $\chi_{z}:=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, and $(\frac{1}{2}\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric.
@@
More generally, the ERM framework also applies to many other problems including regression, functional estimation, matrix completion, shape-restricted regression, and isotonic regression; see [7, 6]. In any case, the inequality (18) plays a key role in upper bounding the mean squared error $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]$; see e.g. [41, eq (2)]. Since $\mathbb{E}[\hat{L}(m)-L(m)]=0$, the expected generalization error equals $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], which is also connected to (18). If we take the expectation of both sides of (18), and upper bound the right side by the Dudley integral using the covering number of $T$, we generally do not get sharp convergence rates of the mean squared error. In fact, even for the simplest case of 1-dimensional parameter estimation, such a naive approach yields a $1/\sqrt{n}$ rate rather than the optimal $1/n$ rate.
@@
The traditional approach for improving the bound is a “slicing” argument that yields local complexity bounds; see for example [41, 44, 35] or [42, Theorem 5.32]. Roughly speaking, if $\|\hat{m}-m\|^{2}=r$, then $\chi_{m-\hat{m}}\leq\sup_{z\in B(r)\cap(T-T)}\chi_{z}$, which is bounded by some function $\omega(r)$ with high probability. The function $\omega(r)$ typically grows slower than $r^{2}$, so $r^{2}\lesssim\frac{1}{\sqrt{n}}\omega(r)$ yields an upper bound on $r$. More precisely, a union bound argument is applied to control the generalization error at different scales of $r$ with high probability. While it may seem that the uniform control by the union bound could lose logarithmic factors, this can actually be avoided by carefully scheduling of the error tolerance at each scale.
@@
Local complexity bounds are usually based on Dudley’ integral, which can be loose for general $T$ when Dudley’ integral is loose. The majorizing measure theorem, though sharp, is hard to evaluate, and is rarely used in the analysis of ERM.
@@ 2.3 Information-theoretic tools
We recall standard information-theoretic measures in [12], using the natural base of the logarithm. The entropy of a discrete random variable $X\sim\mu$ is $H(X):=\sum_{x}\mu(x)\ln\frac{1}{\mu(x)}$. The relative entropy between $\mu$ and $\nu$ on the same measurable space is $D(\mu\|\nu):=\int\ln\frac{d\mu}{d\nu}d\mu$. Mutual information for $(X,Y)\sim P_{XY}$ is defined as $I(X;Y):=D(P_{XY}\|P_{X}\times P_{Y})$. Conditional versions of these quantities are defined by averaging over the random variable which is conditioned on. For example, $D(P_{\hat{Z}|Z}\|\mu|P_{Z}):=\int D(P_{\hat{Z}|Z=z}\|\mu)dP_{Z}(z)$, and $H(Y|X):=\int H(Y|X=x)dP_{X}(x)$, where $H(Y|X=x):=\sum_{y}P_{Y|X=x}(y)\ln\frac{1}{P_{Y|X=x}(y)}$. The lifting argument relies on the method of types, a standard method in information theory [13] and large deviation analysis [14]. The type of a sequence is defined as its empirical distribution. Since not all distributions can be a type for a given length of the sequence, we need the following:
@@
Suppose that $\mu$ is a distribution on a finite set $\mathcal{Z}$. We say $\mu$ is rational if there exists integer $N>0$ such that $N\mu$ is integer, i.e., $N\mu(z)\in\mathbb{Z}$ for any $z\in\mathcal{Z}$.
If $N\nu$ is integer, the type class of $\nu$ is defined as the set of sequences with type $\nu$. The following result is standard and can be found in [13, Lemma 2.6]:
@@
Let $\mathcal{S}_{N}$ be the set of all possible types of $\mathcal{Z}^{N}$, where $\mathcal{Z}$ is a finite set with cardinality $n:=|\mathcal{Z}|$. Then $|\mathcal{S}_{N}|={n+N-1\choose n-1}\leq(N+1)^{n}$. Furthermore, if $\mu$ and $\nu$ are distributions on $\mathcal{Z}$, $N\nu$ is integer, and $\mathcal{C}$ denotes the type class of $\nu$, then
$(N+1)^{-n}\exp(-ND(\nu\|\mu))\leq\mu^{\otimes N}(\mathcal{C})\leq\exp(-ND(\nu\|\mu)).$ (19)
Lemma 1 implies the following result, which can be found in [13, Chapter 2, Problem 3]:
Lemma 2.
Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are finite sets, $P_{XY}$ is a distribution on $\mathcal{X}\times\mathcal{Y}$, and $NP_{XY}$ is integer. Let $x^{N}$ be a sequence of type $P_{X}$, and let $\mathcal{C}\subseteq\mathcal{Y}^{N}$ be the type class of $P_{Y}$. Define
$\mathcal{C}(x^{N}):=\{y^{N}\colon(x^{N},y^{N})\textrm{ has type $P_{XY}$}\}.$ (20)
Then we have
$(N+1)^{-|\mathcal{X}||\mathcal{Y}|}\exp(-NI(X;Y))\leq\frac{|\mathcal{C}(x^{N})|}{|\mathcal{C}|}\leq(N+1)^{|\mathcal{Y}|}\exp(-NI(X;Y)).$ (21)
Note that sequences in $\mathcal{C}(x^{N})$ and $\mathcal{C}$ have the same probability under $P_{Y}^{\otimes N}$. Taking $\mu=\nu=P_{Y}$ in Lemma 1 yields
$\quad(N+1)^{-2|\mathcal{X}||\mathcal{Y}|}\exp(-NI(X;Y))\leq P_{Y}^{\otimes N}(\mathcal{C}(x^{N}))\leq(N+1)^{|\mathcal{Y}|}\exp(-NI(X;Y)).$ (22)
We recall the rate-distortion function in information theory [12], which will be a counterpart of the covering number in Dudley’s integral. Given a probability measure $\mu$ on $T$ and $\sigma>0$, define
Definition 3.
Given $\mu$ on a metric space $(T,d)$, define the rate-distortion function
$R_{\mu}(\sigma^{2}):=\inf_{P_{UZ}}I(U;Z),$ (24)
where the infimum is over probability measure $P_{UZ}$ on $T\times T$ satisfying $P_{Z}=\mu$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$. Define $r_{\mu}(\sigma):=R_{\mu}(\sigma^{2})$, and
[…]
Observe that
$r_{\mu}(\sigma)\leq i_{\mu}(\sigma)\leq r_{\mu}(\sigma/2).$ (26)
[…]
For $Z\sim\mu$, define
$\sigma_{\rm m}(\mu):=\inf_{z\in T}\sqrt{\mathbb{E}[d^{2}(Z,t)]},$ (27)
@@ 3.1 A rate-distortion integral
with ties broken arbitrarily. Assuming measurability of $Z(\omega)$, we have measurability of $d(Z(\omega),t_{j})$ for each $j$, so that $Z_{k}$, and hence $X_{Z_{k}}$, is measurable. Density of $T_{0}$ ensures $\lim_{k\to\infty}Z_{k}=Z$ almost surely, and continuity implies that $X_{Z}=\lim_{k\to\infty}X_{Z_{k}}$ is measurable, being a sequential limit of measurable functions.
@@
Given a stochastic process $(X_{t}(\omega))_{t\in T}$ and distribution $\mu$ on $T$, define the complexity
@@
If $(X_{t})_{t\in T}$ is a centered sub-Gaussian process on a countable set $T$, and $\mu$ is a distribution on $T$ for which $\sigma_{\rm m}:=\sigma_{\rm m}(\mu)<\infty$, then
$\mathsf{w}(\mu)\leq 2K\int_{0}^{\sigma_{\rm m}}\sqrt{i_{\mu}(\sigma)}d\sigma\leq 4K\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (31)
where $K$ is the constant in (6).
Theorem 1 is proved in Section 4 for finite $T$ and extended to countable $T$ in the appendix. It will be shown the appendix that if the right side of (31) is finite, then $\mathbb{E}[|X_{Z}|]<\infty$, so no need to worry about absolute integrability on the left side of (31).
@@
Suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, and let $Y:=(X_{t})\in\mathbb{R}^{|T|}$. Let $Z$ be a random variable on $T$ arbitrarily correlated with $Y$, and set $\mu:=P_{Z}$. Let $K$ be the constant in (6). Then
$\mathbb{E}[X_{Z}]\leq 2K\int_{0}^{\sigma_{\rm m}}\sqrt{i_{\mu}(\sigma)\wedge I(Y;Z)}d\sigma.$ (32)
The proof of Theorem 2 is given in Section 4. Since $\sigma_{\rm m}\lesssim\sup_{t\in T}\|X_{t}\|_{G}$, it is easy to see that Theorem 2 implies (2).
While Theorem 1 does not cover the case of uncountable $T$, it is sufficient for our statistical applications due to the following observation:
@@
If $(X_{t})_{t\in T}$ is a separable, continuous process, then for $Z\sim\mu$,
$\mathbb{E}[X_{Z}]\leq 2K\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(\mu,\mu^{\prime})\leq\epsilon}\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (33)
The proof of Proposition 1 is in the appendix. Naturally, we would like to show that the right side of (33) reduces to the right side of (31) by dominated convergence, but this requires showing that $\sqrt{R_{\mu^{\prime}}(\sigma^{2})}$ can be dominated by an absolutely integrable function, which is challenging in the regime of vanishing $\sigma$ where the rate-distortion function may grow fast. While it is possible to impose on certain growth condition for dominated convergence to apply, we will not dive into those complications. Instead, in many statistical applications, we will see that $R_{\mu^{\prime}}(\sigma^{2})$ can be controlled using the second moments of $\mu$, so the right side of (33) can be directly evaluated using the continuity of the bound in the second moment, without the need to establish the equivalence of the right sides of (33) and (31).